# NPL-MVPS: Neural Point-Light Multi-View Photometric Stereo

### v2

@@ 4.1 Datasets
DiLiGenT-MV. Our main evaluation target is DiLiGenT-MV [24] containing 5 objects with 96 lights in 20 views. Images are $612\times 512$ resolution with objects actually occupying a maximum of $400\times 400\text{px}$. Ground truth meshes, camera intrinsics, extrinsics and normal maps are provided, as well as point light positions and far-field light brightnesses ($\phi$). We note that these brightnesses were measured from the intensity of a flat calibration target roughly positioned at the location of the imaged objects, so intrinsic brightness is recovered by multiplying with inverse distance square. We note that as such calibration data is unavailable, the $\phi$ is expected to be fairly inaccurate and thus it is optimised during training.
@@ 4.2 Hyperparameters and Training
We use a tensorflow port of SIREN for all the experiments. The SDF network is set to 5x512 layers (1.05M parameters) and the albedo network to 3x256 (133K parameters). The BRDF network is set to 3x32 layers (2.5K parameters). We use 64 ray samples in a 100mm ray range and an additional 64 around the average intersection (in a shrinking distance range up to 10mm) and 64 extra samples computed with one step of Newton method (for approximating the 0 of the SDF). For each shadow ray we used 16 samples. We train with batch size of 512 rays for 100 epochs which takes approx. 20h on a NVIDIA TITAN RTX and 17GB of RAM when rendering all 96 lights. Note that the 6 lights 6 views version completes in only 2h.
@@ 4.3 Evaluation protocol
We evaluate our method by computed Chamfer distance (marked as surface error SE) of the reconstructions and the ground truth . This is computed as the average of asymmetric Hausdorf distance from reconstruction to ground truth and the opposite, with the distances computed with with Meshlab. We note that in order to have a fair comparison and not bias the error with unseen bottom of the objects, the bottom 6mm of GTs and all reconstructions are removed. All DiLiGenT-MV objects are aligned to be touching the XY plane so the cropping is straightforward. This cropping also avoids large error at some parts of the bottom of the objects that are occluded by the background (e.g. the feet of Reading, see supplementary). Thus our reported error (in Table 2) is generally lower than the numbers originally reported in other works.
Competing approaches. We compare against DiLiGenT-MV [24], PS-NeRF [44], MVAS [5] Brahimi et. al, [1] RNB  [2] and Supernormal [4]. DiLiGenT-MV [24] and Brahimi et. al, [1] are closed source ones so we use the meshes computed by the original authors; for all other methods we use the meshes we were able to reproduce from their codes. We note that Supernormal [4] offers the best perfomance and by far the least computational cost so it is used for ablating the sparse lights and views scenario as well. Note that Brahimi et. al, [1] was only computed on the sparse scenario therefore that is the only available comparison for them.
@@ 5.1 Synthetic data
As mentioned in Section 4.1 the original DiLiGenT-MV [24] dataset contains several anomalies making hard to correctly ablate the several steps of our method. Thus ablation experiments are performed on Synthetic-DiLiGenT-MV dataset described in Section 4.
We first show that our network can achieve a very low error of 0.05mm when using the ground truth normals which is not the case for the real DiLiGenT-MV dataset as shown in Table 2. We also show that if predicted normals are used the performance is worse: 0.14mm. Using the rendering of intensities only achieves overall accuracy of 0.16mm. In addition, using intensities only significantly outperforms the normals only network on Bear, Cow, Reading but it is inferior on Buddha and Pot2. The reason for such a differing performance is that the presence of strong self reflection effects presents a more difficult task to the rendering network than the normal estimation network. The combined normals and intensity rendering experiment achieves the best accuracy of 0.12mm average error. In addition, on Bear and Cow objects, the accuracy using both losses is better than any of the individual ones (0.17mm, 0.13mm vs 0.11mm for Bear) and (0.10mm, 0.07mm vs 0.04mm for Cow) therefore showing the ability to combine best of both training signals.
@@
Table 2: Results on original dense DiLiGenT-MV [24] benchmark. For all objects we report the Chamfer distance error as well as average error on all objects. Normals are computed using the universal PS method of [12]. We evaluate 3 versions of our method named N, I and N+I namely normal loss only (N) - no rendering, rendering loss only (I) - no normals loss and combined (N+I). We also include comparison with Supernormal (SpN [4]) with ground truth normals (GT N).
@@ 5.2 Real data
Table 3: Results on sparse DiLiGenT-MV [24] benchmark. We include three sparse cases (marked with [views, lights] and compare with SpN [4] which is the best performing dense competitor. We also include the comparison with Brahimi et al [1] which was computed with 6 views and 75 lights which is comparable to our 6 views 30 lights case. We note that we significantly outperform Supernormal [4] (0.38mm vs 0.61mm) on the most sparse setup [6 views, 6 lights] and also confidently outperform Brahimi et. al [1] (0.25mm vs 0.34mm) despite having a significantly smaller number of lights in [6 views, 30 lights] setup. Supernormal [4] matches our performance when more views are used.
Original (dense) DiLiGenT-MV setup. Our method achieves (0.2mm) state of the art results (see Table 2), only matching to the original DiLiGenT-MV benchmark (0.22mm) and Supernormal [4] (0.2mm). Note we significantly outperform most deep learning based competitors (ie. PS-NeRF [44], MVAS [5] and RNB [2]). Also note in PS-NeRF [44] the performance of was compared by including the bottom (invisible) part of the object which gave misleading results of the deep learning method outperforming the classical one proposed originally with the DiLiGenT-MV [24] benchmark. We believe that the classical method performs so well as the objects have relatively simple geometry (especially the Cow where is mostly convex and has the least amount of shadows and self reflection). The reconstructed meshes are shown in Figure 5.
Sparse DiLiGenT-MV setup. In Table 3 provide our results on various sparse setups of DiLiGenT-MV [24] benchmark. In particular, we include three sparse cases (marked with [views, lights]). In all cases we compare with Supernormal [4] which is the best competitor on dense DiLiGenT-MV benchmark and has code available. We also include the comparison with Brahimi et al [1] (code is not available) which was originally computed with 6 views and 75 lights which is comparable to our 6 views 30 lights case. We note that we significantly outperform Supernormal [4] (0.38mm vs 0.61mm) on the most sparse setup [6 views, 6 lights] where the normal estimates are very inaccurate (see Figure 1) and confidently outperform Brahimi et. al [1] (0.25mm vs 0.34mm) despite having a significantly smaller number of lights in [6 views, 30 lights] setup. Supernormal [4] matches (0.34 mm vs 0.35mm) our performance when more views (20) are used.
It is also noteworthy that only using 6 lights and 6 views (around 2% of the total data) only increases the total error from 0.2mm to 0.38mm signifying the need for a more challenging multi-view photometric stereo benchmark where the extra images would be required to achieve sub-millimeter accuracy. Additionally, a more challenging dataset would allow newer approaches to display a larger gain vs the original method, which is still marginally worse than SOTA (0.22mm vs 0.2mm).
@@ 6 Conclusions
Computational efficiency has been beyond the scope of this project, however future work can target that aspect though intergration with the NERFACC [25] framework.
@@ Appendix A Appendix
This document contains a brief discussion of DiLiGenT-MV data anomalies in Section B.
Appendix B DiLiGenT-MV data anomalies

### v3

@@ 4.1 Datasets
DiLiGenT-MV. Our main evaluation is performed on DiLiGenT-MV [25] benchmark containing 5 objects with 96 lights in 20 views. Images are of $612\times 512$ px resolution with objects actually occupying a maximum of $400\times 400\text{px}$. Ground truth meshes, camera intrinsics, extrinsics and normal maps are provided, together with point light positions and far-field light brightnesses ($\phi$). We note that these brightnesses were measured from the intensity of a flat calibration target roughly positioned at the location of the imaged objects, so intrinsic brightness is recovered by multiplying with inverse distance square. We note that as such calibration data is unavailable, the $\phi$ is expected to be fairly inaccurate and thus it is optimised during training.
@@ 4.2 Hyperparameters and Training
We use a tensorflow port of SIREN for all the experiments. The SDF MLP is set to 5x512 layers (1.05M parameters) and the albedo MLP to 3x256 (133K parameters). The BRDF MLP is set to 3x32 layers (2.5K parameters). We use 64 ray samples in a 100mm ray range and an additional 64 around the average intersection (in a shrinking distance range up to 10mm) and 64 extra samples computed with one step of Newton method (for approximating the 0 of the SDF). For each shadow ray we used 16 samples. We train with batch size of 512 rays for 100 epochs which takes approx. 20h on a NVIDIA TITAN RTX and 17GB of RAM when rendering all 96 lights. Note that the 6 lights 6 views version completes in only 2h.
@@ 4.3 Evaluation protocol
We evaluate our method by computed Chamfer distance (marked as surface error SE) of the reconstructions and the ground truth. This is computed as the average of asymmetric Hausdorf distance from reconstruction to ground truth and the opposite, with the distances computed with with Meshlab. We note that in order to have a fair comparison and not bias the error with unseen bottom of the objects, the bottom 6mm of GTs and all reconstructions are removed. All DiLiGenT-MV objects are aligned to be touching the XY plane so the cropping is straightforward. This cropping also avoids large error at some parts of the bottom of the objects that are occluded by the background (e.g. the feet of Reading, see Supplementary Material). Thus our reported error (in Table 2) is generally lower than the numbers originally reported in other works.
Competing approaches. We compare against DiLiGenT-MV [25], PS-NeRF [47], MVAS [5] Brahimi et. al, [1] RNB  [2] and Supernormal [4]. DiLiGenT-MV [25] and Brahimi et. al, [1] are closed source so we use the meshes computed by the original authors; for all other methods we use the meshes reproduced from their original codes. We note that Supernormal [4] offers the best perfomance and by far the least computational cost so it is used for ablating the sparse lights and views scenario as well. Also Brahimi et. al, [1] was only computed on the sparse scenario therefore that is the only available comparison for them.
@@ 5.1 Synthetic data
As mentioned in Section 4.1 the original DiLiGenT-MV [25] dataset contains several anomalies making hard to correctly ablate the several steps of our method. Thus additional ablation experiments are performed on Synthetic-DiLiGenT-MV dataset.
We first show that our network can achieve a very low error of 0.05mm when using the ground truth normals which is not the case for the real DiLiGenT-MV dataset as shown in Table 2. We also show that if predicted normals are used the performance is worse: 0.14mm. Using the rendering of intensities only achieves overall error of 0.12mm. In addition, using the rendering loss on intensities only (I) significantly outperforms the normals loss only (N) experiments on all objects except Buddha. The reason for such a differing performance is that the presence of strong self reflection effects presents a more difficult task to the rendering network than for the normal estimation network. Note that not computing the shadow maps while using rendering only loss (I-S) increases the error from 0.12mm to 0.15mm, with the very concave object Reading being affected the most (0.12m to 0.18mm). The combined normals and intensity rendering experiment achieves the best accuracy of 0.09mm average error.
@@
Table 2: Results on original dense DiLiGenT-MV [25] benchmark. For all objects we report the Chamfer distance error as well as average error on all objects. Normals are computed using the universal PS method of [13]. We evaluate four versions of our method, including using only normal (N) loss, rendering loss with (I) and without shadows (I-S) as in Table 1. We also include comparison with Supernormal (SpN [4]) with ground truth normals (GT N).
@@ 5.2 Real data
Table 3: Results on sparse DiLiGenT-MV [25] benchmark. We include three sparse cases (marked with [views, lights] and compare with Supernormal [4] which is the best performing dense competitor. We also include the comparison with Brahimi et al [1] which was computed with 6 views and 75 lights and is comparable to our 6 views 30 lights case. We note that our method significantly outperform Supernormal [4] (0.38mm vs 0.61mm) on the most sparse setup [6 views, 6 lights] as well as Brahimi et. al [1] (0.25mm vs 0.34mm) in [6 views, 30 lights] setup. Supernormal [4] matches our performance when more views are used.
Original (dense) DiLiGenT-MV setup. Our method achieves (0.2mm) state of the art results (see Table 2), only matching to the original DiLiGenT-MV benchmark (0.22mm) and Supernormal [4] (0.2mm). Note we significantly outperform most deep learning based competitors (ie. PS-NeRF [47], MVAS [5] and RNB [2]). Also note in PS-NeRF [47] the performance of various methods was compared by including the bottom (invisible) part of the object which gave misleading results of the deep learning method outperforming the classical one proposed originally with the DiLiGenT-MV [25] benchmark. We believe that the classical method performs so well as the objects have relatively simple geometry (especially the Cow where is mostly convex and has the least amount of shadows and self reflection). The reconstructed meshes are shown in Figure 5.
Sparse DiLiGenT-MV setup. In Table 3 provide our results on various sparse setups of DiLiGenT-MV [25] benchmark. In particular, we include three sparse cases (marked with [views, lights]). In all cases we compare with Supernormal [4] which is the best competitor on dense DiLiGenT-MV benchmark and has code available. We also include the comparison with Brahimi et al [1] (code is not available) which was originally computed with 6 views and 75 lights which is comparable to our 6 views 30 lights case. Our method significantly outperforms Supernormal [4] (0.38mm vs 0.61mm) on the most sparse setup [6 views, 6 lights] where the normal estimates are very inaccurate (see Figure 1) and significantly outperforms Brahimi et. al [1] (0.25mm vs 0.34mm) despite having a significantly smaller number of lights in [6 views, 30 lights] setup. Supernormal [4] matches (0.34 mm vs 0.35mm) our performance when more views (20) are used. Visualisation of sparse results can be found in the supplementary material.
It is also noteworthy that only using 6 lights and 6 views (around 2% of the total data) only increases the total error from 0.2mm to 0.38mm signifying the need for a more challenging multi-view photometric stereo benchmark.
@@ 6 Conclusions
Finally, it is important to note that improving computational efficiency has been beyond the scope of this project, however future work can improve it significantly by following strategies proposed by Supernormal [4] and integrating with the NERFACC [26] framework.
@@ Appendix A Appendix
This appendix contains qualitative results on DiLiGenT-MV [25] benchmark in Section B and a brief discussion of DiLiGenT-MV data anomalies in Section C.
Appendix B Qualitative results on sparse DiLiGenT-MV [25]
Qualitative results on sparse DiLiGenT-MV [25] benchmark are shown in Figure 6. We include all 3 cases namely [6 views, 6 lights], [6 views, 30 lights] and [20 views, 6 lights]. For us, we include best version (N+I) and compare with SpN [4], and Brahimi et al [1], which was computed with 6 views and 75 lights which is comparable to our 6 views 30 lights case. We observe that our methods achieves consistent low error in most regions of most objects and it is thus the overall best competitor. It is notable, that Brahimi et al [1] does not model cast shadows and thus achives high error in concavities like between the legs of Bear and the inside of Reading.
Figure 6: Qualitative results on the sparse version of real DiLiGenT-MV [25] benchmark. The square bracket for each case denote number of [views, lights]. For each mesh vertex, the minimum distance to the GT mesh is shown with the error bars set to red corresponding to 1mm.
Appendix C DiLiGenT-MV data anomalies